OpenStudy (anonymous):

Three points X = (3, 0, 1) Y = (4, 5, 0) Z = (−3, −8, −1) Are they collinear?

ganeshie8 (ganeshie8):

what do you know about cross product of two vectors ?

OpenStudy (anonymous):

Collinear points have same slope

OpenStudy (anonymous):

Can u find the points of them

OpenStudy (mayankdevnani):

@hopelovelift do you know the property of collinear points ?

OpenStudy (anonymous):

I mean slope

ganeshie8 (ganeshie8):

interesting, how can we find slope in 3 dimensions ?

OpenStudy (mayankdevnani):

yeah,too interesting

OpenStudy (anonymous):

Okay i'll show my working. So to find line/vector I did this XY= Y-X and YZ=Z-Y Keep in mind that I gave you points not three vectors :P Then I assumed that Collinear points are Parallel. XY=kYZ Then I found out that it is not parallel therefore not collinear.

OpenStudy (mayankdevnani):

@Raja99 can you tell me how to find slope in 3 dimensions?

OpenStudy (anonymous):

points wont be parallel.. online lines

OpenStudy (anonymous):

only*

OpenStudy (anonymous):

Also I have another equation if you guys don't mind. Let A = (2, 1, 3) and B = (−1, −1, −1). Find the vector parametric equation of the line through A and B. Using your equation determine a point C, distinct from A and B such that A, B, C are collinear. I Just want to know how to solve it because I missed the lectures on parametric equation...

OpenStudy (mayankdevnani):

@hopelovelift i tell you a shortcut. Shortcut:- Find discriminant of all the points and equate to zero. Because AREA OF COLLINEAR POINTS IS ZERO. 

OpenStudy (anonymous):

no idea.. but can we look in 2d like xy..xz..yz

OpenStudy (anonymous):

What is the graph showing? Are you telling me its collinear or not?

OpenStudy (mayankdevnani):

see,if area of all the points become zero,then we can definitely say that these points are collinear and if they don't get zero,so they are not collinear points. And to find area ,use discriminant.

OpenStudy (mayankdevnani):

understood? @hopelovelift

OpenStudy (anonymous):

Okay thanks.

OpenStudy (mayankdevnani):

welcome :)

OpenStudy (anonymous):

Do you also know how to solve the question above?

OpenStudy (mayankdevnani):

use discriminant,i said above

OpenStudy (anonymous):

To find the parametric equation?

OpenStudy (mayankdevnani):

no....

OpenStudy (mayankdevnani):

https://www.youtube.com/watch?v=21LWuY8i6Hw this will help you so much.... Using this method,if this is equal to 0,then these are collinear points

OpenStudy (anonymous):

Okay thanks

OpenStudy (mayankdevnani):

welcome :)